OpenStudy (anonymous):

If the probability of seeing moose in a day in a certain region of Alaska is 0.74, what is the probability of not seeing any moose in a day there?

OpenStudy (anonymous):

probibility is out of 1

OpenStudy (anonymous):

so 1- chance of seeing= chance of not seeing

OpenStudy (anonymous):

i have 4 options. .74 .26 1.26 -.74

OpenStudy (anonymous):

what do you think?

OpenStudy (anonymous):

In a more formal way : p(x) = probability of event x. Given Omega the universe of possible outcome, P(Omega) = 1. Given P(Moose) = 0.74, the complementary event "Not seeing moose" is by definition : P(Not moose) = P(Omega) - P(Moose). P(Not moose) = 1 - 0.74. P(Not Moose) = 0.26

OpenStudy (anonymous):

.-74